# Gaitsgory’s central functor and the Arkhipov–Bezrukavnikov equivalence in mixed characteristic

### v3

@@ Abstract.
We show that the nearby cycles functor for the $p$-adic Hecke stack at parahoric level is perverse t-exact, by developing a theory of Wakimoto filtrations at Iwahori level, and that it lifts to the ${\mathbb{E}}_{1}$-center. We apply these tools to construct the Arkhipov–Bezrukavnikov functor for $p$-adic affine flag varieties at Iwahori level, and prove that it is an equivalence for type $A$ groups.
@@ 1. Introduction
Let $F$ be a $p$-adic field with ring of integers $O$ and residue field $k$, and let $G$ be a connected reductive $F$-group with a parahoric $O$-model ${\mathcal{G}}$. The first goal of this paper is to define a $p$-adic analogue of the Gaitsgory central functor [Gai01] sending perverse sheaves on the Hecke stack ${\operatorname{Hk}}_{G,C}$ over the completed algebraic closure $C$ of $F$ to central perverse sheaves on the Hecke stack ${\operatorname{Hk}}_{{\mathcal{G}},\bar{k}}$. The second goal of the paper is, when ${\mathcal{G}}={\mathcal{I}}$ is Iwahori, to also define a $p$-adic analogue of the Arkhipov–Bezrukavnikov functor [AB09] relating coherent sheaves on the dual Springer variety $\hat{{\mathcal{N}}}_{\mathrm{Spr}}$ to constructible étale sheaves on ${\operatorname{Hk}}_{{\mathcal{I}},\bar{k}}$. During the introduction, we will assume for simplicity that $G$ is split and that the coefficients of our sheaves equal $\bar{\mathbb{Q}}_{\ell}$, as these hypothesis hold for most of the paper. We begin by recalling some of the representation-theoretic aspects at the level of Grothendieck groups.
@@ 1.1. Hecke algebras
where $\bar{\mathbb{Q}}_{\ell}$ is regarded as a ${\mathbb{Z}}[q^{\pm 1/2}]$-algebra by mapping $q$ to the cardinality of $k$ (and thus choosing a square root of this integer in $\bar{\mathbb{Q}}_{\ell}$). Let ${\mathbb{H}}_{\mathrm{f}}\subset{\mathbb{H}}$ be the finite Hecke algebra associated with the finite Weyl group and the set of simple reflections. The antispherical module is defined as
@@
where $T_{w}\in{\mathbb{H}}_{\mathrm{f}}$ acts by multiplication by ($1)^{\ell({w})}q^{1/2}$ on ${\mathbb{Z}}[q^{\pm 1/2}]^{\mathrm{sgn}}$. We also define the antispherical module for ${\mathcal{H}}$ to be ${\mathcal{M}}^{\mathrm{as}}:={\mathbb{M}}\otimes_{{\mathbb{Z}}[q^{\pm 1/2}]}\bar{\mathbb{Q}}_{\ell}$.
@@
The Steinberg variety is defined as $\hat{\mathrm{St}}:=\hat{{\mathcal{N}}}_{\mathrm{Spr}}\times_{\hat{{\mathcal{N}}}}\hat{{\mathcal{N}}}_{\mathrm{Spr}}$. Kazhdan-Lusztig [KL87] showed that the affine Hecke algebra is isomorphic to the Grothendieck group $K_{0}([\hat{G}\times{\mathbb{G}}_{m}\backslash\hat{\mathrm{St}}])$, where the latter has an algebra structure induced by convolution. In particular, the antispherical module ${\mathbb{M}}^{\mathrm{as}}$ is identified with $K_{0}([\hat{G}\times{\mathbb{G}}_{m}\backslash\hat{{\mathcal{N}}}_{\mathrm{Spr}}])$. If we forget the ${\mathbb{G}}_{m}$-equivariance, then we recover both the Iwahori-Hecke algebra ${\mathcal{H}}$ and its anti-spherical module ${\mathcal{M}}^{\mathrm{as}}$.
On the other hand, it follows from the work of Iwahori–Matsumoto [IM65] that the Iwahori-Hecke algebra coincides with $K_{0}({\mathscr{P}}({\operatorname{Hk}}_{{\mathcal{I}}}))$ where ${\mathscr{P}}$ denotes the category of perverse sheaves on the Hecke stack ${\operatorname{Hk}}_{{\mathcal{I}}}=L^{+}{\mathcal{I}}\backslash LG/L^{+}{\mathcal{I}}$. The natural action of ${\mathcal{H}}$ on ${\mathcal{M}}^{\mathrm{as}}$ induces a surjective map ${\mathcal{H}}\rightarrow{\mathcal{M}}^{\mathrm{as}}$ with kernel generated by the Kazhdan-Lusztig basis elements indexed by the $w\in W$, which are not minimal in their left $W_{\mathrm{fin}}$-coset. This leads us to consider the antispherical category of perverse sheaves ${\mathscr{P}}_{\mathrm{as}}({\operatorname{Hk}}_{{\mathcal{I}}})$ given as the Serre quotient by the ${\operatorname{IC}}$-sheaves index by those $w$. Another approach to realize ${\mathcal{M}}^{\mathrm{as}}$ is as the $I$-invariants of the compact induction to $G(F)$ of a generic character $\chi$ of the unipotent radical $I_{u}^{\mathrm{op}}$ of the opposite Iwahori subgroup. This is the so-called Iwahori–Whittaker model and its categorification plays an important role later in our arguments.
@@ 1.2. The central functor
Let us explain a bit of the notions and ingredients that go into the above theorem. Our treatment of the centrality of ${\mathscr{Z}}$ is to our best knowledge the only one that uses the machinery of stable $\infty$-categories, which entails additional higher coherent homotopies. An important technical tool is the notion of an abstract six-functor formalism in the sense of Mann [Man22a, Man22b], which allows us to work at the level of the category of correspondences. Once we are there, we perform the usual fusion trick of looking at the disjoint locus of $(\mathrm{Spd}(O_{C}))^{2}$ and conclude the desired monoidality via full faithfulness of pullback away from the diagonal for those sheaves which are perverse over $(\mathrm{Spd}C)^{2}$. To obtain this full faithfulness, we apply a certain calculation of nearby cycles of kimberlites from [GL22].
Trying to prove perversity of ${\mathscr{Z}}(V)$ was the genesis of this project. Contrary to the function field case, we cannot rely on Artin vanishing to provide us with this crucial fact. Instead, we first consider a Iwahori ${\mathcal{I}}$ and look at the Wakimoto functor ${\mathscr{I}}\colon{\operatorname{Rep}}(\hat{T})\to{\mathscr{P}}({\operatorname{Hk}}_{\mathcal{I}})$ following [AB09], but defined instead at the level of complexes as in [AR]. The centrality of ${\mathscr{Z}}(V)$ implies that it lies in the full subcategory generated by the essential image of ${\mathscr{I}}$ under extensions. Each graded piece can then be recovered by invoking geometric Satake and a certain orthogonality with respect to the constant terms $\mathrm{CT}_{B^{\mathrm{op}}}$, which proves perversity and the existence of a Wakimoto filtration all at once. This differs considerably from the strategy in [AB09], which exploits both perversity and convolution-exactness (known in equicharacteristic by Artin vanishing). Perversity in the general parahoric case can be deduced from the Iwahori one, based on a suggestion of Achar, which we learned from Cass–Scholbach–van den Hoven [CvdHS24], who adapted our argument for Iwahori models to their setting. From the Wakimoto filtration, we can also deduce the convolution-exactness and unipotency of the monodromy operator induced by the Galois group (note that $\mathrm{Sat}(V)$ descends to $\mathrm{Spd}(F)$ with trivial inertia action). In the meantime, the Wakimoto filtration has been decisively used in [GL22] to give a new proof of unibranchness (i.e., topological normality) of local models in complete generality.
@@ 1.3. The AB functor and Iwahori–Whittaker sheaves
If $G$ is of type $A$, then ${\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)$ is tilting.
Notice that in [AB09] there is no assumption on the type of $G$. Unfortunately, we are missing a crucial ingredient replacing the ${\mathbb{G}}_{m,k}$-action given by rotating the uniformizer, which is impossible in the $p$-adic setting. Let us explain how the proof works, so as to better understand the impact of this gap. First of all, the tilting property propagates under convolution and can be verified on adjoint quotients, so we may assume $V$ is either minuscule or quasi-minuscule. In the minuscule case, all the weights are comprised in a finite Weyl orbit, so one can easily verify the given property. In the quasi-minuscule case, we must still handle the (co)restriction of ${\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)$ to the weight 00 Iwahori–Whittaker cell. Here, the vanishing can be achieved by calculating the alternating sum of the $\mathrm{Ext}$ groups via an argument on Grothendieck groups and finally bound a $\mathrm{Hom}({\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(1),{\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V))$ accordingly. Now, this bound is achieved in [AB09] via the theory of the regular quotient described below together with the fact that the monodromy operator is defined for every sheaf in ${\mathscr{P}}({\operatorname{Hk}}_{{\mathcal{I}}})$ as it is induced by the ${\mathbb{G}}_{m,k}$-action given by rotation. For type $A$ groups, we can realise every representation up to central isogeny as a direct summand of a minuscule representation, so we still get the tilting property.
The last step in proving that ${\mathscr{F}}_{{\mathcal{I}}{\mathcal{W}}}$ is an equivalence (now, necessarily only in type $A$) revolves around the regular orbit ${\mathcal{O}}_{\mathrm{r}}\subset\hat{{\mathcal{N}}}$ inside the nilpotent cone. The Springer resolution is an isomorphism above this $\hat{G}$-orbit, and hence we should find a category of étale sheaves that plays a similar role. For this, we look at the Serre quotient ${\mathscr{P}}_{0}({\operatorname{Hk}}_{{\mathcal{I}}})$ of perverse sheaves on the Hecke stack obtained by modding out IC sheaves with positive-dimensional support.
@@
If $G$ is of type $A$, then there is a symmetric monoidal equivalence ${\mathscr{P}}_{0}({\operatorname{Hk}}_{{\mathcal{I}}})\xrightarrow{\sim}\mathrm{Rep}(Z_{\hat{G}}(n_{0}))$, where $n_{0}$ is a regular nilpotent element.
Together with the tilting property, this result is key in order to prove the Arkhipov–Bezrukavnikov equivalence, as it induces certain injections of $\mathrm{Hom}$ maps. Let us remark that the most delicate point in the above theorem consists in showing that $n_{0}$ is regular. For this, we use the theory of weights by descending ${\mathscr{Z}}(V)$ to a mixed sheaf and calculating its monodromy filtration. In [AB09], one applies Gabber’s local weight-monodromy theorem, see [BB93], stating that the weight filtration equals the monodromy filtration, and then calculates the former via the affine Hecke algebra. This is not available for our nearby cycles, unless $\mu$ is minuscule, by work of Hansen–Zavyalov [HZ23] combined with the representability theorem in [AGLR22]. Again, we can only easily reproduce this argument for type $A$ groups. It would be possible to adapt a different argument due to Bezrukavnikov–Riche–Rider [BRR20], but this would lead us into some detours that seem unnecessary, as we do not have the tilting property for other groups.
Let us finish by stating the second main result of this paper, i.e. the AB equivalence for $p$-adic groups of type $A$:
@@
If $G$ is of type $A$, then the functor ${\mathscr{F}}_{{\mathcal{I}}{\mathcal{W}}}\colon\mathrm{Perf}([\hat{G}\backslash\hat{{\mathcal{N}}}_{\mathrm{Spr}}])\to{\mathscr{D}}_{\mathrm{cons}}({\operatorname{Hk}}_{{\mathcal{I}}{\mathcal{W}}})$ is an equivalence.
We strongly believe that this result must also hold for general split connected reductive groups $G$. Yun–Zhu have announced in conference talks regarding work of preparation of Hemo–Zhu, see also [Zhu20], a proof of the full Bezrukavnikov equivalence [Bez16] for $p$-adic groups, that builds on a colimit presentation in terms of double quotients of parahoric jet groups due to Tao–Travkin [TT20]. Recently, Bando [Ban23] also gave a distinct proof of the Bezrukavnikov equivalence for $p$-adic groups by comparing constructible-étale sheaves in equi- and mixed characteristic via an ingenious geometric construction. However, these previous methods do not yield concrete knowledge about the central functor, whereas our paper places ${\mathscr{Z}}$ right at the center of it all. We also think that our functor ${\mathscr{Z}}$ will naturally appear in the picture if one studies étale sheaves on $p$-adic Hecke stacks, see, e.g., the unibranchness theorem of [GL22], and thus it must play a role in comparing the Zhu [Zhu20] and the Fargues–Scholze [FS21] variants of a categorical $p$-adic local Langlands correspondence. A natural task for the future will be to explain if and how all of the previous approaches fit together, namely by comparing a priori different central functors.

### v4

@@ Abstract.
We show that the nearby cycles functor for the $p$-adic Hecke stack at parahoric level is perverse t-exact, by developing a theory of Wakimoto filtrations at Iwahori level, and that it lifts to the ${\mathbb{E}}_{1}$-center. We apply these tools to construct the Arkhipov–Bezrukavnikov functor for $p$-adic affine flag varieties at Iwahori level, and prove that it is an equivalence for all classical groups and also exceptional groups of type $E_{6}$ and $E_{7}$.
@@ 1. Introduction
Let $F$ be a $p$-adic field with ring of integers $O$ and residue field $k$, and let $G$ be a connected reductive $F$-group with a parahoric $O$-model ${\mathcal{G}}$. The first goal of this paper is to define a $p$-adic analogue of the Gaitsgory central functor [Gai01] sending perverse sheaves on the Hecke stack ${\operatorname{Hk}}_{G,C}$ over the completed algebraic closure $C$ of $F$ to central perverse sheaves on the Hecke stack ${\operatorname{Hk}}_{{\mathcal{G}},\bar{k}}$. The second goal of the paper is, when ${\mathcal{G}}={\mathcal{I}}$ is Iwahori, to also define a $p$-adic analogue of the Arkhipov–Bezrukavnikov functor [AB09] relating coherent sheaves on the dual Springer variety $\hat{{\mathcal{N}}}_{\mathrm{Spr}}$ to constructible étale sheaves on ${\operatorname{Hk}}_{{\mathcal{I}},\bar{k}}$. During the introduction, we will assume for simplicity that $G$ is split and that the coefficients of our sheaves equal $\bar{\mathbb{Q}}_{\ell}$, as these hypotheses hold for most of the paper. We begin by recalling some of the representation-theoretic aspects at the level of Grothendieck groups.
@@ 1.1. Hecke algebras
where $\bar{\mathbb{Q}}_{\ell}$ is regarded as a ${\mathbb{Z}}[q^{\pm 1/2}]$-algebra by mapping $q$ to the cardinality of $k$ (and thus choosing a square root of this integer in $\bar{\mathbb{Q}}_{\ell}$). Let ${\mathbb{H}}_{\mathrm{f}}\subset{\mathbb{H}}$ be the finite Hecke algebra associated with the finite Weyl group and the set of simple reflections. The anti-spherical module is defined as
@@
where $T_{w}\in{\mathbb{H}}_{\mathrm{f}}$ acts by multiplication by ($1)^{\ell({w})}q^{1/2}$ on ${\mathbb{Z}}[q^{\pm 1/2}]^{\mathrm{sgn}}$. We also define the anti-spherical module for ${\mathcal{H}}$ to be ${\mathcal{M}}^{\mathrm{as}}:={\mathbb{M}}\otimes_{{\mathbb{Z}}[q^{\pm 1/2}]}\bar{\mathbb{Q}}_{\ell}$.
@@
The Steinberg variety is defined as $\hat{\mathrm{St}}:=\hat{{\mathcal{N}}}_{\mathrm{Spr}}\times_{\hat{{\mathcal{N}}}}\hat{{\mathcal{N}}}_{\mathrm{Spr}}$. Kazhdan-Lusztig [KL87] showed that the affine Hecke algebra is isomorphic to the Grothendieck group $K_{0}([\hat{G}\times{\mathbb{G}}_{m}\backslash\hat{\mathrm{St}}])$, where the latter has an algebra structure induced by convolution. In particular, the anti-spherical module ${\mathbb{M}}^{\mathrm{as}}$ is identified with $K_{0}([\hat{G}\times{\mathbb{G}}_{m}\backslash\hat{{\mathcal{N}}}_{\mathrm{Spr}}])$. If we forget the ${\mathbb{G}}_{m}$-equivariance, then we recover both the Iwahori-Hecke algebra ${\mathcal{H}}$ and its anti-spherical module ${\mathcal{M}}^{\mathrm{as}}$.
On the other hand, it follows from the work of Iwahori–Matsumoto [IM65] that the Iwahori-Hecke algebra coincides with $K_{0}({\mathscr{P}}({\operatorname{Hk}}_{{\mathcal{I}}}))$ where ${\mathscr{P}}$ denotes the category of perverse sheaves on the Hecke stack ${\operatorname{Hk}}_{{\mathcal{I}}}=L^{+}{\mathcal{I}}\backslash LG/L^{+}{\mathcal{I}}$. The natural action of ${\mathcal{H}}$ on ${\mathcal{M}}^{\mathrm{as}}$ induces a surjective map ${\mathcal{H}}\rightarrow{\mathcal{M}}^{\mathrm{as}}$ with kernel generated by the Kazhdan-Lusztig basis elements indexed by the $w\in W$, which are not minimal in their left $W_{\mathrm{fin}}$-coset. This leads us to consider the anti-spherical category of perverse sheaves ${\mathscr{P}}_{\mathrm{as}}({\operatorname{Hk}}_{{\mathcal{I}}})$ given as the Serre quotient by the ${\operatorname{IC}}$-sheaves index by those $w$. Another approach to realize ${\mathcal{M}}^{\mathrm{as}}$ is as the $I$-invariants of the compact induction to $G(F)$ of a generic character $\chi$ of the unipotent radical $I_{u}^{\mathrm{op}}$ of the opposite Iwahori subgroup. This is the so-called Iwahori–Whittaker model and its categorification plays an important role later in our arguments.
@@ 1.2. The central functor
Let us explain a bit of the notions and ingredients that go into the above theorem. Our treatment of the centrality of ${\mathscr{Z}}$ is to our best knowledge the only one that uses the machinery of stable $\infty$-categories, which entails additional higher coherent homotopies. An important technical tool is the notion of an abstract six-functor formalism in the sense of Mann [Man22a, Man22b], which allows us to work at the level of the category of correspondences. Once we are there, we perform the usual fusion trick of looking at the disjoint locus of $(\mathrm{Spd}(O_{C}))^{2}$ and conclude the desired monoidality via full faithfulness of pullback away from the diagonal for those sheaves which are perverse over $(\mathrm{Spd}C)^{2}$. To obtain this full faithfulness, we apply a certain calculation of nearby cycles of kimberlites from [GL24].
Trying to prove perversity of ${\mathscr{Z}}(V)$ was the genesis of this project. Contrary to the function field case, we cannot rely on Artin vanishing to provide us with this crucial fact. Instead, we first consider an Iwahori ${\mathcal{I}}$ and look at the Wakimoto functor ${\mathscr{I}}\colon{\operatorname{Rep}}(\hat{T})\to{\mathscr{P}}({\operatorname{Hk}}_{\mathcal{I}})$ following [AB09], but defined instead at the level of complexes as in [AR]. The centrality of ${\mathscr{Z}}(V)$ implies that it lies in the full subcategory generated by the essential image of ${\mathscr{I}}$ under extensions. Each graded piece can then be recovered by invoking geometric Satake and a certain orthogonality with respect to the constant terms $\mathrm{CT}_{B^{\mathrm{op}}}$, which proves perversity and the existence of a Wakimoto filtration all at once. This differs considerably from the strategy in [AB09], which exploits both perversity and convolution-exactness (known in equicharacteristic by Artin vanishing). Perversity in the general parahoric case can be deduced from the Iwahori one, based on a suggestion of Achar, which we learned from Cass–Scholbach–van den Hove [CvdHS24], who adapted our argument for Iwahori models to their setting. From the Wakimoto filtration, we can also deduce the convolution-exactness and unipotency of the monodromy operator induced by the Galois group (note that $\mathrm{Sat}(V)$ descends to $\mathrm{Spd}(F)$ with trivial inertia action). In the meantime, the Wakimoto filtration has been decisively used in [GL24] to give a new proof of unibranchness (i.e., topological normality) of local models in complete generality.
@@ 1.3. The AB functor and Iwahori–Whittaker sheaves
If $G$ has enough minuscules, then ${\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)$ is tilting.
Notice that in [AB09] there is no assumption on $G$. Unfortunately, we are missing a crucial ingredient replacing the $\mathbb{G}_{m,k}$-action given by rotating the uniformizer, which is impossible in the $p$-adic setting. Contrary to what was asserted in an initial version of this paper, the impact of this gap is very mild, and we get the result for all groups with enough minuscules, see Definition 7.8, a class comprised of all classical groups, and also exceptional ones of types $E_{6}$ and $E_{7}$. This oversight on our side was brought to our attention by the work of Dhillon–Taylor [DT25].
Let us explain how the proof works, so the reader can better grasp the gap above. First of all, the tilting property propagates under convolution and can be verified on adjoint quotients, so we may assume $V$ is either minuscule or quasi-minuscule by a lemma of Ngô–Polo [NP01]. In the minuscule case, all the weights are comprised in a finite Weyl orbit, so one can easily verify the given property. In the quasi-minuscule case, we must still handle the (co)restriction of ${\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)$ to the weight 00 Iwahori–Whittaker cell. Here, the vanishing can be achieved by calculating the alternating sum of the $\mathrm{Ext}$ groups via an argument on Grothendieck groups and finally bound the dimension of the $\bar{{\mathbb{Q}}}_{\ell}$-vector space $\mathrm{Hom}({\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(1),{\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V))$ accordingly. Now, this bound is achieved in [AB09] via the theory of the regular quotient described below together with the fact that the monodromy operator is defined for every sheaf in ${\mathscr{P}}({\operatorname{Hk}}_{{\mathcal{I}}})$ as it is induced by the ${\mathbb{G}}_{m,k}$-action given by rotation. For groups with enough minuscules, we can realise every representation up to central isogeny as a direct summand of a tensor product of minuscule representations, so we still get the tilting property.
The last step in proving that ${\mathscr{F}}_{{\mathcal{I}}{\mathcal{W}}}$ is an equivalence (now, necessarily assuming enough minuscules) revolves around the regular orbit ${\mathcal{O}}_{\mathrm{r}}\subset\hat{{\mathcal{N}}}$ inside the nilpotent cone. The Springer resolution is an isomorphism above this $\hat{G}$-orbit, and hence we should find a category of étale sheaves that plays a similar role. For this, we look at the Serre quotient ${\mathscr{P}}_{0}({\operatorname{Hk}}_{{\mathcal{I}}})$ of perverse sheaves on the Hecke stack obtained by modding out IC sheaves with positive-dimensional support.
@@
If $G$ has enough minuscules, then there is a symmetric monoidal equivalence ${\mathscr{P}}_{0}({\operatorname{Hk}}_{{\mathcal{I}}})\xrightarrow{\sim}\mathrm{Rep}(Z_{\hat{G}}(n_{0}))$, where $n_{0}$ is a regular nilpotent element.
Together with the tilting property, this result is key in order to prove the Arkhipov–Bezrukavnikov equivalence, as it induces certain injections of $\mathrm{Hom}$ maps. Let us remark that the most delicate point in the above theorem consists in showing that $n_{0}$ is regular. For this, we use the theory of weights by descending ${\mathscr{Z}}(V)$ to a mixed sheaf and calculating its monodromy filtration. In [AB09], one applies Gabber’s local weight-monodromy theorem, see [BB93], stating that the weight filtration equals the monodromy filtration, and then calculates the former via the affine Hecke algebra. This is not available for our nearby cycles, unless $\mu$ is minuscule, by work of Hansen–Zavyalov [HZ23] combined with the representability theorem in [AGLR22]. Again, we can only easily reproduce this argument for groups with enough minuscules. It would be possible to adapt a different argument due to Bezrukavnikov–Riche–Rider [BRR20], but this would lead us into some detours that seem unnecessary, as we do not have the tilting property for other groups.
Let us finish by stating the second main result of this paper, i.e. the AB equivalence for $p$-adic groups with enough minuscules.
@@
If $G$ has enough minuscules, then the functor ${\mathscr{F}}_{{\mathcal{I}}{\mathcal{W}}}\colon\mathrm{Perf}([\hat{G}\backslash\hat{{\mathcal{N}}}_{\mathrm{Spr}}])\to{\mathscr{D}}_{\mathrm{cons}}({\operatorname{Hk}}_{{\mathcal{I}}{\mathcal{W}}})$ is an equivalence.
Let us remark that in Corollary 9.2 we identify the anti-spherical Serre quotient $\mathscr{P}_{\mathrm{as}}(\mathrm{Hk}_{\mathcal{I}})$ with the category $\mathscr{P}(\mathrm{Hk}_{\mathcal{IW}})$ of Iwahori–Whittaker perverse sheaves, so upon passing to Grothendieck groups, the equivalence above really recovers the classical setup explained in the beginning of this introduction. We strongly believe that the theorem above must also hold for general split connected reductive groups $G$. Yun–Zhu have announced in conference talks regarding work of preparation of Hemo–Zhu, see also [Zhu20], a proof of the full Bezrukavnikov equivalence [Bez16] for $p$-adic groups, that builds on a colimit presentation in terms of double quotients of parahoric jet groups due to Tao–Travkin [TT20]. Recently, Bando [Ban23] also gave a distinct proof of the Bezrukavnikov equivalence for $p$-adic groups by comparing constructible-étale sheaves in equi- and mixed characteristic via an ingenious geometric construction. However, these previous methods do not yield concrete knowledge about the central functor, whereas our paper places ${\mathscr{Z}}$ right at the center of it all. We also think that our functor ${\mathscr{Z}}$ will naturally appear in the picture if one studies étale sheaves on $p$-adic Hecke stacks, see, e.g., the unibranchness theorem of [GL24], and thus it must play a role in comparing the Zhu [Zhu20] and the Fargues–Scholze [FS21] variants of a categorical $p$-adic local Langlands correspondence. A natural task for the future will be to explain if and how all of the previous approaches fit together, namely by comparing a priori different central functors.